OpenStudy (jt24680):

john can cover 84 cm with each stride he takes while running. If he runs 2.5 km, how many strides will he need to take

OpenStudy (ambuj):

1km=100000cm 1cm=1/100000 km 84cm=84/100000 km 84/100000 km=1stride 2.5 km = 100000*2.5/84 strides

OpenStudy (ambuj):

2976 is ans